# Should You Use Your Large Language Model to Explore or Exploit?

### v2

@@ Abstract
We evaluate the ability of the current generation of large language models (LLMs) to help a decision-making agent facing an exploration-exploitation tradeoff. We use LLMs to explore and exploit in silos in various (contextual) bandit tasks. We find that while the current LLMs often struggle to exploit, in-context mitigations may be used to substantially improve performance for small-scale tasks. However even then, LLMs perform worse than a simple linear regression. On the other hand, we find that LLMs do help at exploring large action spaces with inherent semantics, by suggesting suitable candidates to explore.
@@ 1 Introduction
There’s been significant interest in the machine learning community to apply recent advances in generative AI and large language models (LLMs) to solve important decision-making problems. Early work in this direction has already produced impressive agentic behavior in both virtual (e.g.,  Wang et al., 2023; OpenAI, 2025) and physical-world environments (e.g.,  Black et al., 2024).
Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making the best decision given the current data) and exploration (trying new options for long-term benefit). Balancing the two has led to an enormous literature (e.g.,  see books Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
@@
A recent line of work (e.g.,  Krishnamurthy et al., 2024; Nie et al., 2024) evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  specifying the problem description, parameters, and history in the LLM prompt. Focused on simple tasks in reinforcement learning (RL), these results are mixed. Both papers show that LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context. However, such statistics do not exist beyond simple decision-making problems, e.g.,  for contextual bandits. RL-specific pre-training or fine-tuning (on data from algorithmic baselines on similar problem instances) tends to work well (see Related Work), but may be prohibitive due to cost or insufficient training data. Besides, using a commonly available frontier LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
@@
Motivated by these observations, we study the ability of LLMs to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent. We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff. We experiment with a variety of LLMs: Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, as well as DeepSeek-R1, a reasoning model.  [^2] Several other LLMs did not work for our purposes: Llama-3.2-3B and Llama-3-8B would not follow instructions, as well as DeepSeek-R1-Distill-Llama-70B (a reasoning model); Qwen-2.5-Math-70B (another reasoning model that we tried) has a short context window. We restrict our GPT access to Gpt-4, Gpt-4o, and Gpt-3.5, as newer variants such as Gpt-4.1 and Gpt-4.5 are primarily optimized for efficiency (e.g. latency, cost, coding) rather than broad general reasoning, and Gpt-5 uses a reasoning backbone as a routing mechanism across sub-models.
@@
In Section 2, we evaluate LLMs as exploitation oracles for contextual bandits. Given a history of (context, action, reward) tuples, the LLM is tasked with identifying the best action to take given a new context. Our results here are mixed. We show that LLMs can effectively exploit in-context for small-sized problems, but their performance degrades when the problem becomes moderately sized. We find that in-context summary techniques are useful for improving performance, but LLMs with these mitigations still perform worse than a simple linear regression baseline.
@@
In Section 3, we evaluate LLMs as exploration oracles which suggest candidate actions by discretizing a large action space. To do so, we introduce a text-based multi-armed bandit task, where actions correspond to free-text answers to an open-ended question, and rewards are driven by the distance from some preselected answer in an embedding space (the latter is computed exogenously). Given the high dimensionality of the action space, traditional discretization approaches typically used in continuous bandit problems are inapplicable. However, an LLM can generate a small set of candidate actions which can then be used to instantiate an off-the-shelf bandit algorithm. We experiment with several prompting strategies, and find that they all lead to relatively good exploration. Finally, we repeat our experiments on a larger-scale bandit task based on paper titles and abstracts from arXiv (where the goal is to find a suitable title for a given abstract), with similar findings.
@@
A broader literature on in-context learning (starting from Brown et al., 2020) aims to solve various tasks by providing all relevant information in the LLM prompt. The work on exemplar selection (selecting examples and other information to present in-context) (e.g.,  Khalifa et al., 2023; Zhang et al., 2022; Xiong et al., 2023; Tonglet et al., 2023) is relevant to our exploitation experiments.
@@
Background on multi-armed bandits (MAB). We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract the explore-exploit tradeoff, see Slivkins (2019); Lattimore and Szepesvári (2020) for background. In MAB, there are $T$ rounds and $K$ arms. In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some sub-Gaussian reward distribution for this arm. The reward distribution, and particularly its mean $\mu(a_{t})$, are unknown to the algorithm. In CB, the learner additionally observes a context $z_{t}$ before each round $t$, and the expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm. The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
An “exploitation oracle” (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader. Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification (Slivkins, 2019, Ch.8). Designing CB exploitation oracles for their own sake, a.k.a. offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford, 2009; Dudík et al., 2014). Usage of an LLM as an “exploration oracle” is closely related to the literature on dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011), as we elaborate in Section 3.
@@ 2 LLMs as exploitation oracles
Figure 1: MAB exploit puzzle for Gpt-4 (left), Gpt-4 with CoT (middle), and Gpt-3.5 with CoT (right), all with ”buttons” prompt. The following conventions apply to all figures in this section. Each line corresponds to a particular value of #rounds $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis. The shaded band around the line represents a $95\%$ confidence interval. The dashed line is the number of tasks (“runs”) with empirical gap $\leq\epsilon$; the resp. Y-scale is on the right.
We evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes. We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB). In a CB exploit task, an LLM is given a history consisting of context-arm-reward tuples, and is instructed to take the best arm given the current history and the current context. A MAB exploit task is the same, but without contexts. These tasks are generated from some parameterized distributions called exploit puzzles.
As our results here are largely negative, we focus on relatively small tasks. We use synthetic data, to better control task size/difficulty and trace the “boundary of tractability”. We find that LLMs are consistently outperformed by a simple linear regression baseline, even on moderately-sized tasks. Due to computational constraints, our experiments focused on Gpt-4, Gpt-4o, and Gpt-3.5, with additional LLMs evaluated as a robustness check in Appendix A.4, with similar findings.
MAB exploit puzzles. Our MAB-based experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for why the current LLMs fail to solve MAB tasks in-context when presented with raw (non-summarized) history, as first observed by Krishnamurthy et al. (2024); Nie et al. (2024). Following these two papers, we try two prompts: one in which arms correspond to pushing different colored buttons and one where they correspond to showing different advertisements to users. The LLM is asked to choose the arm with the highest empirical reward in the next round. We also try chain-of-thought (CoT) prompts, for the total of 4 prompt designs: $\left\{\,\text{buttons, adverts}\,\right\}\times\left\{\,\text{CoT, no-CoT}\,\right\}$. See Appendix A for more details on our experimental setup.
@@
Given a (realized) exploit task, we measure its difficulty via empirical gap $\epsilon$: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$. Intuitively, puzzle difficulty decreases with $\epsilon$ (as one needs to distinguish the top two arms). Empirical gap mirrors the “gap” between the top two expected rewards, a standard notion of difficulty in MAB. Let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$.
@@
We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer”: an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$. We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level. Hence, we plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against empirical gap $\epsilon$.
We find that Gpt-4 and Gpt-3.5 do not perform well on these MAB exploit puzzles, see Figure 1. Performance tends to degrade (1) as the history length $T$ increases and (2) as the empirical gap decreases. (Our intuition for the former is that the LLM has more data points to process in context, and for the latter is that the bandit instance becomes more difficult.) While Gpt-4 generally performs much better than Gpt-3.5, we found that prompting the LLM to use chain-of-thought (CoT) reasoning provided a slight boost for Gpt-3.5, while hurting the performance of Gpt-4.333From CoT logs, we found that Gpt-4 was doing incorrect reasoning in a non-trivial fraction of runs. We found that performance was similar across all of our prompt designs (see Appendix A for more plots).
Figure 2: Gpt-4 succeeds on a small CB exploit puzzle (left), but fails on a slightly larger one (right).
@@
CB exploit puzzles and mitigations. While the history in $K$-armed bandits can be summarized using $2K$ numbers (for each arm, the average reward and #plays) such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB. [^4] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
@@
We focus on linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) parameters $\theta^{*}_{a}\in\mathbb{R}^{d}$.  [^5] Linear CB are well-studied, starting from Li et al. (2010); Chu et al. (2011); Abbasi-Yadkori et al. (2011). (We try non-linear CB, too.)
@@
We consider a CB exploit puzzle parameterized by #arms $K$, dimension $d$, and history size $T$. The tasks are constructed as follows. We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-0.25,0.25]$ independently and u.a.r. for each arm $a$. Given context $z\in\mathbb{R}^{d}$, expected reward for arm $a$ is $\mu(z,a)=\langle z,\theta_{a}\rangle+\gamma_{a}$. We generate a history of $T$ rounds $t\in[T]$. Contexts $z_{t}$ are sampled independently and u.a.r. from $1,1]^{d}$. For simplicity, the history contains rewards of all arms $a$ in each round $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian. Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action for round $T+1$ which appears best. This gives one exploit task. We generate $N$ tasks for the same $K,T,d$.
@@
Given an exploit task, we now define a “correct answer” as an arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$.  [^6] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context. $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks. Likewise, the task difficulty is not easily defined in terms of the realized rewards. Instead, we focus on the effective gap: the difference in expected reward between the best and second-best arm given the current context $z=z_{T+1}$. That is, the difference between the largest and second-largest number among $\mu(a,z)$, $a\in[K]$. Intuitively, smaller effective gap corresponds to increased difficulty.
We are interested in how $\mathrm{FracCorrect}$ varies with effective gap. In each plot, we fix the number of arms ($K$) and dimension ($d$), and let $S(\epsilon,T)$ be the set of all tasks with given $K,d,T$ and effective gap at most $\epsilon$. We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
First, we find that Gpt-4 obtains near-perfect performance on a “small” puzzle with $K=2$ arms, context dimension $d=1$, and history size $T=50$ (Figure 2). However, its performance degrades as the problem size increases: see Figure 2 for $K=d=2$ and $T=100$. Moreover, limited prompt size may prevent processing larger histories. [^7] E.g.,  our LLM access points bottomed out at $T\approx\text{100-200}$ for Gpt-4 and $T\approx\text{1000-2000}$ for Gpt-4o.
Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially. Gpt-4 without CoT (left) and Gpt-4 with CoT (right). Note that providing the full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.
Motivated by these observations, we implement several natural mitigations inspired by the literature on exemplar selection for in-context learning (discussed in Related Work).
1. $k$-nearest: Among the observed contexts, consider the distinct $k$ contexts closest to $z_{T+1}$, according to the $\ell_{2}$ metric. Limit the history reported in the prompt to (the rounds with) these $k$ contexts.
@@
We do not explain the “mitigation strategy” in the prompt: we present the reported context-arm-reward tuples as if it were the entire history, without mentioning clustering or averaging.
@@
Figure 3 visualizes the performance of these mitigations on a slightly larger (but still relatively small) puzzle with $K=d=2$ and $T=4000$. We use Gpt-4 with and without CoT. We compare the mitigations against linear regression baseline (which is effectively an ”upper bound”, as the underlying CB instance is linear). Without CoT prompting, we find that all three mitigations achieve $\mathrm{FracCorrect}$ around $80\%-90\%$, although this dips to around $60\%-85\%$ when using CoT. In addition to (potentially) improving performance, mitigations can also offer a practical way to solve decision-making tasks using LLMs when the history is large; when $T=4,000$, our prompt vastly exceeds the context window of all models we had access to.
@@
Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
@@
However, current LLMs struggle to exploit on even moderately-sized problems, even with these mitigations. In Figure 4, we plot the performance of Gpt-4o with mitigations under various hyperparameters. While $k$-means (left) is almost as bad as random guessing, the $k$-nearest mitigations (center) achieve about $50\%$ $\mathrm{FracCorrect}$, and the $k$-means, $k^{\prime}$-nearest mitigations (right) obtain approximately $60\%$ $\mathrm{FracCorrect}$. Both our $k$-nearest and $k$-means, $k^{\prime}$-nearest mitigations significantly out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
Figure 5: Left: Performance of Deepseek-R1 on our numerical CB puzzle. Right: Gpt-4o on the text-based CB exploit puzzle. Some mitigations help, but are outperformed by linear regression.
Reasoning models are LLMs specifically trained to follow multi-step chains of thought to solve complex problems that require intermediate reasoning. We ran one such model, DeepSeek-R1-Distill-Qwen-32B (henceforth DeepSeek-R1), a 32B parameter reasoning model distilled from Qwen-2.5. Our results are shown in Figure 5 (left). As before, we find that DeepSeek-R1 fails to match the performance of the baseline, even when using the mitigations. High inference costs and small context windows prevented us from running larger reasoning models.
@@
CB exploit puzzles (text-based & non-linear). As a robustness check, we repeat our CB experiments on a text-based exploit puzzle. In this puzzle, contexts are items in a room (e.g. animals, objects on a table), and actions have an associated semantic meaning (e.g. eat the food item, leave the room). Rewards are still presented numerically, and are non-linear functions of both the context and action. See Appendix A for full details on our experimental setup.
@@
Figure 5 (right) shows the performance of Gpt-4o (with mitigations) on this puzzle. While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline.

### v3

@@ Abstract
We evaluate the ability of the current generation of large language models (LLMs) to help a decision-making agent facing an exploration-exploitation tradeoff. While previous work has largely study the ability of LLMs to solve combined exploration-exploitation tasks, we take a more systematic approach and use LLMs to explore and exploit in silos in various (contextual) bandit tasks. We find that reasoning models show the most promise for solving exploitation tasks, although they are still too expensive or too slow to be used in many practical settings. Motivated by this, we study tool use and in-context summarization using non-reasoning models. We find that these mitigations may be used to substantially improve performance on medium-difficulty tasks, however even then, all LLMs we study perform worse than a simple linear regression, even in non-linear settings. On the other hand, we find that LLMs do help at exploring large action spaces with inherent semantics, by suggesting suitable candidates to explore.
@@ 1 Introduction
The machine learning community is increasingly interested to apply advances in generative AI and large language models (LLMs) to decision-making problems. Early work in this direction has already produced impressive agentic behavior in both virtual (e.g.,  Wang et al., 2023; OpenAI, 2025) and physical-world environments (e.g.,  Black et al., 2024). Beyond generalization (needed for supervised learning), decision-making under uncertainty requires two additional capabilities: exploitation (making best decisions given the current data) and exploration (for long-term benefit). Balancing the two has led to a large literature (e.g.,  Slivkins, 2019; Lattimore and Szepesvári, 2020; Agarwal et al., 2020).
@@
A recent line of work (e.g.,  Krishnamurthy et al., 2024; Nie et al., 2024) evaluates the ability of LLMs to balance exploration and exploitation entirely in-context, i.e.,  specifying the problem description, parameters, and history in the LLM prompt. Focused on simple tasks in reinforcement learning (RL), these results are mixed. LLMs fail to solve these tasks adequately out-of-the-box, but they can be prompted to do so by providing succinct summary statistics in-context. However, such statistics do not exist beyond simple RL problems, e.g.,  for contextual bandits. RL-specific pre-training or fine-tuning (on data from algorithmic baselines on similar problem instances) tends to work well (see Related Work), but may be prohibitive due to cost or insufficient training data. Besides, using a commonly available LLM would leverage its “generalist” intelligence and may be much easier in terms of logistics and required expertise.
@@
Motivated by these observations, we study the ability of LLMs to explore and exploit in-context in silos, with an eye towards leveraging a pre-trained LLM (and the inductive bias therein) as a part of a larger decision-making agent. We focus on (contextual) bandits, as a standard abstraction for the explore-exploit tradeoff. We experiment with many LLMs: Gpt-5.2, Gpt-5-mini, Gpt-5-nano, Gpt-4, Gpt-4o, Gpt-3.5, Qwen-2.5, Gemma-3, Mistral-7B, and DeepSeek-R1, a reasoning model.  [^2] Some other LLMs did not work for our purposes: Llama-3.2-3B would not follow instructions. Two reasoning models, DeepSeek-R1-Distill-Llama-70B and Qwen-2.5-Math-70B have a context window that is too short.
@@
In Section 2, we evaluate LLMs as exploitation oracles for contextual bandits. Given a history of (context, action, reward) tuples, the LLM is tasked with identifying the best action to take given a new context. Our results here are mixed. We show that LLMs can effectively exploit in-context for small-sized tasks but their performance degrades when the tasks become moderately sized. We find that frontier reasoning models outperform non-reasoning models (fixing model size and provider), but tend to be slow and/or expensive. We investigate to mitigations: in-context summary techniques and tool use (namely, a Python code interpreter). We find that both mitigations improve performance in a linear task, but on on-linear tasks LLMs with these mitigations still perform worse than a simple linear regression baseline.
@@
In Section 3, we evaluate LLMs as an exploration oracle which suggests a small set of candidate actions by discretizing a large action space. (This set can then be used to instantiate an off-the-shelf bandit algorithm.) On various text-based multi-armed bandit tasks that we design, we find this approach far superior to a non-LLM baseline. Our first task is based on the MovieLens dataset (Harper and Konstan, 2015), where actions are movie recommendations, rewards are random draws parameterized by ground-truth movie ratings, and the baseline is a hand-crafted discretization based on movie genre. Second, we design a larger-scale bandit task based on paper titles/abstracts from arXiv: find a suitable title for a given abstract. Given the high dimensionality of the action space, discretization approaches from continuous bandit problems are inapplicable. Several prompting strategies all lead to relatively good exploration compared to natural baselines. Our third task is based on open-ended “philosophical” questions and contrarian answers.
@@
A broader literature on in-context learning (starting from Brown et al. (2020)) aims to solve various tasks by providing all relevant information in the LLM prompt. The work on exemplar selection (selecting examples and other information to present in-context) (e.g.,  Khalifa et al., 2023; Zhang et al., 2022; Xiong et al., 2023; Tonglet et al., 2023) is relevant to our exploitation experiments.
@@
Background on multi-armed bandits (MAB). We consider tasks based on MAB and contextual bandits (CB), well-studied special cases of RL that abstract explore-exploit tradeoff (Slivkins, 2019; Lattimore and Szepesvári, 2020). In MAB, there are $T$ rounds and $K$ arms. In each round $t\in[T]$, the learner chooses an action (arm) $a_{t}\in[K]$ and observes reward $r_{t}$ drawn from some (fixed, unknown) sub-Gaussian reward distribution $D(a_{t})$ with mean $\mu(a_{t})$. In CB, the learner also observes a context $z_{t}$ before each round $t$. The expected reward $\mu(z_{t},a_{t})$ depends on both the context and the arm. The learner’s goal is to balance exploration and exploitation to maximize cumulative reward.
An “exploitation oracle” (which optimizes for the current round given the history) naturally plugs into standard bandit algorithms such as Epsilon-Greedy, Explore-then-Commit, and Follow-The-Perturbed-Leader. Typical implementations in CB involve model-based (e.g.,  linear) regression or cost-sensitive classification (Slivkins, 2019, Ch.8). Designing CB exploitation oracles for their own sake, a.k.a. offline policy optimization, is well-studied (starting from, e.g.,  Beygelzimer and Langford (2009); Dudík et al. (2014)). Usage of an LLM as an “exploration oracle” is closely related to the literature on dynamic pricing and Lipschitz bandits (e.g.,  Kleinberg and Leighton, 2003; Kleinberg et al., 2008, 2019; Bubeck et al., 2011), as we elaborate in Section 3.
@@ 2 LLMs as exploitation oracles
Figure 1: MAB exploit puzzle for Gpt-3.5 (left), Gpt-4 (middle), and Gpt-5-nano (right), all with the “buttons” prompt and temperature 11. Results for Gpt-5-mini are not visualized, as $\mathrm{FracCorrect}(\epsilon,T)=1$ for all $\epsilon$ and $T$. The following conventions apply to all figures in this section. Each line corresponds to a particular value of #rounds $T$ and plots $\mathrm{FracCorrect}(\epsilon,T)$ against empirical gap $\epsilon$ on the X-axis. The shaded band around the line represents a $95\%$ confidence interval. The dashed line is the number of tasks (“runs”) with empirical gap $\leq\epsilon$; the resp. Y-scale is on the right.
We evaluate the ability of LLMs to exploit in decision-making tasks with statistical uncertainty on the outcomes. We present LLMs with in-context exploit tasks inspired by multi-armed bandits (MAB) and contextual bandits (CB). In a CB exploit task, an LLM is given a history consisting of context-arm-reward tuples, and is instructed to take the best arm given the current history and the current context. A MAB exploit task is the same, but without contexts. These tasks are generated from some parameterized distributions called exploit puzzles. Due to computational constraints, our experiments focus mostly on Gpt model family, with additional LLMs evaluated as a robustness check in Appendix B.4, with similar findings. Across all experiments, varying the model temperature did not significantly affect performance. We include results for temperature 0 and 1.
MAB exploit puzzles. We begin by testing the Gpt model family on simple MAB exploit puzzles. We find that performance significantly improves as the models progress from Gpt-3.5 to Gpt-4 to Gpt-5. In fact, Gpt-5-mini achieves perfect performance. Our experiments on Gpt-4 and Gpt-3.5 provide a partial explanation for the observation in (Krishnamurthy et al., 2024; Nie et al., 2024), that these models fail to solve end-to-end MAB tasks in-context when presented with raw (non-summarized) history.
Following these two papers, we try two prompts, where arms correspond, resp., to pushing different colored buttons and to showing different ads to users. The LLM is asked to choose the arm with the highest empirical reward in the next round. We also try chain-of-thought (CoT) prompts, for the total of 4 prompt designs: $\left\{\,\text{buttons, adverts}\,\right\}\times\left\{\,\text{CoT, no-CoT}\,\right\}$. See Appendix B.1 for more details on this setup.
@@
Figure 2: Gpt-5 (left) and Gpt-4 (right) model families on a numerical CB puzzle (all with temperature 11). Gpt-5 models generally perform better than Gpt-4 models of similar size, although letting Gpt-4 and Gpt-4o use a code interpreter improves performance.
@@
Given a (realized) exploit task, we measure its difficulty via empirical gap $\epsilon$: the difference between the largest and second-largest average reward $\bar{r}_{t}(a):=\frac{1}{T}\sum_{t\in[T]}r_{t}(a)$ among all arms $a$. Intuitively, puzzle difficulty decreases with $\epsilon$ (as it gets harder to distinguish the top two arms). Empirical gap mirrors the “gap” between the top two expected rewards, a standard notion of difficulty in MAB.
@@
We measure an LLM’s performance over a given set $S$ of tasks as the fraction of tasks for which the LLM returns a “correct answer”: an arm with the largest empirical reward; denoted by $\mathrm{FracCorrect}(S)$. We are interested in how $\mathrm{FracCorrect}$ varies depending on the difficulty level. Hence, let $S(\epsilon,T)$ be the set of all tasks with empirical gap at most $\epsilon$ and history size $T$. We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
We find that while Gpt-4 and Gpt-3.5 do not perform well, Gpt-5 models do, see Figure 1. For Gpt-4 and Gpt-3.5, performance tends to degrade as (1) history size $T$ increases, with more data points to process in-context and (2) empirical gap decreases, making the bandit instance more difficult.
@@
CB exploit puzzles. While the history in $K$-armed bandits can be summarized with $2K$ numbers (for each arm, the average reward and #plays), such succinct summary statistics may not be readily available (or even exist) in more complicated decision-making tasks such as CB. [^3] This consideration also motivates MAB exploit puzzles with raw (non-summarized) history, as a simpler special case of the general scenario when succinct summarization is unavailable.
@@
We begin with linear CB, where the expected reward of each arm $a$ is linear in the context $z\in\mathbb{R}^{d}$: $\mu(z,a)=\langle z,\theta^{*}_{a}\rangle$ for some fixed (but unknown) vector $\theta^{*}_{a}\in\mathbb{R}^{d}$.  [^4] Linear CB are well-studied, starting from Li et al. (2010); Chu et al. (2011); Abbasi-Yadkori et al. (2011). (We turn to non-linear CB later in the section.)
@@
We consider a CB exploit puzzle parameterized by #arms $K$, dimension $d$, and history size $T$. The tasks are constructed as follows. We sample parameters $\theta_{a}\in[-1,1]^{d}$ and $\gamma_{a}\in[-\frac{1}{4},\frac{1}{4}]$ independently u.a.r. for each arm $a$. Given context $z\in\mathbb{R}^{d}$, expected reward for arm $a$ is $\mu(z,a)=\langle z,\theta_{a}\rangle+\gamma_{a}$. We generate a history of $T$ rounds $t\in[T]$. Contexts $z_{t}$ are sampled independently u.a.r. from $1,1]^{d}$. For simplicity, the history contains rewards of all arms $a$ in all rounds $t$, where the reward equals $\mu(z_{t},a)$ plus an independent unit-variance Gaussian. Given the history and a new context $z_{T+1}$ (drawn in the same way), the LLM is asked to select the action for round $T+1$ which appears best. This gives one exploit task. We generate $N$ tasks for the same $K,T,d$.
@@
Given an exploit task, we now define a “correct answer” as arm $a$ which maximizes expected reward $\mu(z_{T+1},a)$. [^5] Note that it is unclear how to define an ”empirically best arm” given a CB history and the current context. $\mathrm{FracCorrect}(S)$ is the fraction of correct answers in a given set $S$ of tasks. Task difficulty is also not easily defined in terms of realized rewards. Instead, we focus on effective gap $\epsilon$: the difference in expected reward $\mu(a,z)$ between the best and second-best arm given the current context $z=z_{T+1}$. Intuitively, smaller $\epsilon$ corresponds to increased difficulty.
We study how $\mathrm{FracCorrect}$ varies with effective gap. In each plot, we let $S(\epsilon,T)$ be the set of all tasks with given $K,d,T$ and effective gap at most $\epsilon$. We plot $\mathrm{FracCorrect}(\epsilon,T):=\mathrm{FracCorrect}\left(\,S(\epsilon,T)\,\right)$ against $\epsilon$.
Figure 3: CB exploit puzzle with $d=K=2$ and $T=4000$: mitigations help substantially. Gpt-4 without CoT (left) and Gpt-4 with CoT (right). (Full history with this $T$ vastly exceeds the context window for Gpt-4, Gpt-4o, and Gpt-3.5.)
Figure 2 considers a setting with $d=K=2$ for the Gpt-5 family (left) and the Gpt-4 family (right). We find that larger models generally perform better than their smaller, distilled counterparts and, conditioned on model size, Gpt-5 models outperform Gpt-4 models. We also plot a linear regression baseline (blue dashed line), which effectively serves as an upper-bound on the performance of any model, due to the underlying linear structure of our CB puzzle.
While not pictured, all models tend to perform strictly better on smaller problem instances. For example, Gpt-4 performed near-perfectly when $d=1$, $K=2$, and $T=50$. However, performance degrades as the problem size grows (e.g., see Gpt-4o in Figure 4). Moreover, limited prompt size may prevent processing larger histories. [^6] E.g.,  our LLM access points bottomed out at $T\approx\text{100-200}$ for Gpt-4 and $T\approx\text{1000-2000}$ for Gpt-4o.
Motivated by these observations, we implement two types of mitigations: tool use (providing the model with access to a Python interpreter to execute code in) and in-context summarization (inspired by the literature on exemplar selection for in-context learning, see Related Work). Evaluating these mitigations, we focus on the Gpt-4 family as (1) at the time of writing Gpt-5 models are not configured for tool use and (2) we found the Gpt-5 models to be either prohibitively expensive (Gpt-5.2) or prohibitively slow (Gpt-5-mini and Gpt-5-nano) on larger-scale experiments.
See Figure 2 (right) for the performance of Gpt-4 and Gpt-4o with access to a code interpreter (CI). Both see significantly improved performance with CI, with Gpt-4 even surpassing Gpt-5.2. However, enabling the CI made Gpt-4 and Gpt-4o much slower and much more expensive. [^7] E.g., our Gpt-4 experiment with CI in Figure 2 cost approximately $\mathdollar 70$ and took 6+ hours to run. Our Gpt-4 experiment in Figure 2 without CI cost a few dollars and took under 5 minutes.
We consider the following summarization mitigations:
1. $k$-nearest: Among the observed contexts, pick distinct $k$ contexts closest to $z_{T+1}$, according to the $\ell_{2}$ metric. Limit the reported history to (the rounds with) these $k$ contexts.
@@
We do not explain the “mitigation strategy” in the prompt: we present the context-arm-reward tuples as if it were the entire history without mentioning clustering or averaging.
@@
Figure 3 shows the performance of our mitigations on a slightly larger puzzle with $K=d=2$, $T=4000$. We use Gpt-4 with and without chain-of-thought (CoT) prompting. As before, we compare against a linear regression baseline. Without CoT, all three mitigations achieve $\mathrm{FracCorrect}$ $\approx 80\%-90\%$, although this dips to $\approx 60\%-85\%$ with CoT. Performance aside, the mitigations offer a practical way to handle large histories; e.g., when $T=4,000$, our prompt exceeds the context window of all models we had access to.
@@
Figure 4: CB exploit puzzle with $d=K=5$ and $T=1000$: in-context mitigations perform badly, but (mostly) much better than the no-mitigation baseline. Gpt-4o without CoT.
@@
However, in-context mitigations can only help so much. In Figure 4, we plot their performance under various hyperparameters, on Gpt-4o. While $k$-means (left) is almost as bad as random guessing, $k$-nearest (center) achieves $\approx 50\%$ $\mathrm{FracCorrect}$, and the $k$-means, $k^{\prime}$-nearest mitigation (right) obtains $\approx 60\%$. The latter two out-perform unmitigated Gpt-4o, but fall significantly short of the linear baseline.
@@
CB exploit puzzles (text-based & non-linear). As a robustness check, we repeat our CB experiments on a text-based exploit puzzle. In this puzzle, contexts are items in a room (e.g. animals, objects on a table), and actions have an associated semantic meaning (e.g. eat the food item, leave the room). Rewards are still presented numerically, and are non-linear functions of both the context and action. See Appendix B for full details on our experimental setup.
@@
Figure 5 shows the performance of Gpt-5.2 (left) and Gpt-4o (right) with mitigations on this puzzle. While the reward function is non-linear (and thus the linear baseline only achieves  $70\%$ $\mathrm{FracCorrect}$), we find that all configurations are still significantly out-performed by the linear baseline. Interestingly, Gpt-4o with code interpreter access (orange) performs the worst out of all configurations.